If you travel at 120km/h, how far can you go in 3hours and 15 minutes?
Answer by Saleye(10)

You've given V & t
Simply substitute in the formula...
Given:
V=120 Km/hr
t=3 hrs + (15min./60 min/hr)<= express time with similar unit...
t=13/4 hrs= 3.25 hrs.
Required:
How far(S),distance
=> S=Vt=120 Km/hr * 13/4 hr=390 Km.